# Canonical Policy: Learning Canonical 3D Representation for Equivariant Policy

### v1

@@ V Simulation Experiments
In this section, we provide a comprehensive evaluation of canonical policy in simulation environments. We first present an experiment to validate the $\mathrm{SE}(3)$ equivariance property of canonical policy. We then conduct benchmark comparisons with several representative point cloud-based baselines across a diverse set of manipulation tasks. Finally, we perform ablation studies to highlight the contributions of key components in our method.
@@ V-A Experiment Setup
DP3 [3] is a diffusion-based policy that directly processes 3D point clouds using a simple MLP encoder.
iDP3 [6] builds upon DP3 by replacing the MLP encoder with a multi-layer 1D convolutional encoder. This architecture is designed to extract multi-resolution features from point clouds, capturing more detailed and hierarchical point cloud information.
EquiBot [8] is a SO(3)-equivariant policy that uses the same equivariant encoder as canonical policy, built on the Vector Neuron framework [45]. It modifies the diffusion U-Net backbone using equivariant linear, pooling, and normalization layers. To conform with the Vector Neuron framework, EquiBot reformulates low-dimensional states and actions into vector and scalar representations.
Figure 5: Visualization of the 12 simulated manipulation tasks used in the simulation benchmark. Subfigures (a)–(h) are from MimicGen [51], (i) is from Diffusion Policy [1], and (j)–(l) are from EquiBot [8]. In each subfigure, the left image shows the initial state, and the right image shows the goal state of the task.
@@
MimicGen Tasks [51] are visualized in Fig. 5(a)–(h). These tasks, such as coffee making and table organization, simulate realistic household manipulation scenarios. Point cloud observations are obtained from four camera views. To ensure computational efficiency without compromising performance, all point clouds are uniformly downsampled to 256 points using farthest point sampling [46].
Push T Task is shown in Fig. 5(i). Originally a 2D task from Diffusion Policy [1], we adapt it to 3D by extending each of the 9 keypoints with a zero z-coordinate, forming a sparse 3D point cloud. The goal is to push a T-shaped block to a placed T-shaped target.
Bimanual Tasks are depicted in Fig. 5(j)–(l). These dual-arm manipulation tasks are adapted from the released EquiBot dataset [8], which provides 1024 point clouds as observation.
@@
In canonical policy, we employ absolute control for the MimicGen and Push T tasks, and relative control for the bimanual tasks. The difference between the two control modes lies in the choice of action space used during the SE(3) inverse and forward transformations. See Section IV-C for more details.
@@ V-B Validation of SE(3) Equivariance
Figure 6: Canonical pose visualization across tasks. In-distribution and out-of-distribution observations are aligned to a shared canonical frame via estimated SE(3) inverse transforms.
Figure 7: SE(3) equivariance validation setup. Policies are trained on original data and tested on randomly SE(3)-transformed observations. Actions are mapped back to the original frame for consistency evaluation.
Figure 8: Quantivitily results of different policies when evaluate with in-distribution data and out-of-distribution data.
@@
As illustrated in Fig. 6, we visualize three MimicGen tasks: Stack D1, Hammer Cleanup D1, and Coffee D2, to demonstrate the SE(3)-equivariance property of canonical policy. As discussed in Section IV-A, the left column of Fig. 6 presents point cloud observations from the training distribution, which are considered in-distribution data and denoted as $\mathcal{D}$. Next, random SE(3) transformations are applied to generate out-of-distribution observations $\mathcal{D}^{\mathrm{rd}}$, shown in the middle column. Finally, by applying data canonicalization pipeline described in Section IV-D, we first decenter each transformed point cloud by subtracting its mean, then estimate the associated rotation matrix using an SO(3)-equivariant network (Equation (10)), and finally apply a derotation to align the point cloud to a unique canonical pose $\mathcal{D}^{\mathrm{cn}}$, as visualized in the right column of Fig. 6.
To quantitatively evaluate the SE(3)-equivariance of canonical policy, we design the experimental pipeline depicted in Fig. 7. For each compared policy, we train the model using data $\mathbf{o}\in\mathcal{D}$. During evaluation, we apply random SE(3) transformations to generate $\mathbf{o}^{\mathrm{rd}}\in\mathcal{D}^{\mathrm{rd}}$, similar to the visual setup in Fig. 6.
In Fig. 7(a), a standard policy directly receives the transformed observation $\mathbf{o}^{\mathrm{rd}}$ and predicts the corresponding denoised action $\mathbf{a}^{\mathrm{rd}}$. The same SE(3) transformation that was applied to the observation is then applied inversely to $\mathbf{a}^{\mathrm{rd}}$ to recover the original action $\mathbf{a}$, which is then executed by the robot. In contrast, as shown in Fig. 7(b), canonical policy first estimates the canonical representation ${\mathbf{o}}^{\mathrm{cn}}$ from the transformed observation $\mathbf{o}^{\mathrm{rd}}$. The action ${\mathbf{a}}^{\mathrm{rd}}$ is then predicted from the canonical observation ${\mathbf{o}}^{\mathrm{rd}}$ using the same policy architecture as in the standard pipeline. Finally, a forward SE(3) transformation is applied to map ${\mathbf{a}}^{\mathrm{rd}}$ back to $\mathbf{a}^{\mathrm{rd}}$ in the transformed space, which is then transformed back to the original action $\mathbf{a}$ using the known inverse.
Following this evaluation pipeline, we report results on the Push T and the three MimicGen tasks: Stack D1, Hammer Cleanup D1, and Coffee D2. The results are summarized in Fig. 8. In the figure, red curves indicate the case where the policy is both trained and evaluated on the same distribution $\mathcal{D}$; under this condition, all policies perform well (except EquiBot on MimicGen tasks, which is further analyzed in the ablation study). However, when training is conducted on $\mathcal{D}$ and evaluation on $\mathcal{D}^{\mathrm{rd}}$, only canonical policy achieves stable performance. This strongly validates the benefit of canonical pose representation in enabling robust generalization under arbitrary SE(3) transformations.
@@ V-C Benchmark with Point Cloud Policies
As discussed in Section IV-D, we provide two versions of canonical policy: one with translation invariance and full $\mathrm{SO}(3)$ equivariance, referred to as CP-$\mathrm{SO}(3)$; and another with translation invariance and in-plane $\mathrm{SO}(2)$ equivariance, referred to as CP-$\mathrm{SO}(2)$. We will use these abbreviations throughout the following experimental section.
Table I summarizes the performance of various policies across 12 tasks. Specifically, for the 8 MimicGen tasks, we generate 200 demonstrations [9] and train each method for 250 epochs. For the Push T task, which also contains 200 demonstrations [1], each method is trained for 1000 epochs. For the bimanual tasks, each with 50 demonstrations [8], training is conducted for 2000 epochs per method.
All experiments are conducted using three random environment seeds: 42, 43, and 44. For each seed, we evaluate the policies over the last 10 epochs and report the average. For MimicGen and Push T, each rollout is conducted across 50 different environment initializations. For the three bimanual tasks, each rollout consists of 10 single trials. Thus, the mean score reported in Table I for each task and each method is the average over 30 evaluation results. The best result for each task is shown in bold, and the second-best is underlined.
As shown in Table I, both CP-SO3 and CP-SO2 outperform the baseline point cloud policies (DP3, iDP3, and EquiBot) on all 8 MimicGen tasks and the 3D Push T task. Notably, CP-SO2 achieves the highest success rates in 6 out of the 9 single-arm tasks, supporting our hypothesis that $\mathrm{SO}(2)$ equivariance is particularly well-suited for scenarios where the $z$-axis remains consistently upward.
In the three bimanual tasks, all policies perform similarly in Object Covering and Box Closing, achieving success rates close to 100%. However, for Cloth Folding, CP-SO3 and CP-SO2 achieve the second and third best performances, respectively, following EquiBot.
@@
Interestingly, while EquiBot achieves the best performance across all three bimanual tasks and ranks third on the 3D Push T task, it performs poorly on the MimicGen tasks. We hypothesize that this discrepancy arises from the point cloud encoder used in EquiBot [3], which will further explored in our ablation study on point cloud encoders.
TABLE I: Task success rates (%) across different policies and tasks.
@@
iDP3 $40\pm 6$ $45\pm 7$ $76\pm 1$ $91\pm 6$ $99\pm 1$ $100\pm 0$
EquiBot $1\pm 1$ $5\pm 2$ $77\pm 4$ $98\pm 2$ $99\pm 1$ $100\pm 0$
CP-SO3 $37\pm 6$ $68\pm 6$ $84\pm 3$ $97\pm 6$ $99\pm 1$ $99\pm 6$
CP-SO2 $51\pm 6$ $67\pm 6$ $89\pm 3$ $95\pm 6$ $99\pm 1$ $100\pm 0$
We report the mean success rate, computed as the average over the final 10 checkpoints, each evaluated in 50 environment initializations. Results are further averaged over 3 training seeds, yielding 150 evaluations per policy. Bold indicates the best performance; underline denotes the second-best. “CP-SO3” refers to canonical policy with SO(3) rotation equivariance, and “CP-SO2” denotes the variant with SO(2) equivariance.
@@ V-D Ablation Study
We conducted comprehensive ablation studies focusing on three aspects: (1) different modules within canonical policy, (2) various point cloud encoders, and (3) different action heads. In all experiments, success rates are computed as the mean over the final 10 checkpoints, each evaluated under 50 environment initializations with random seed 42.
As shown in Table II, the abbreviations “Aggre.”, “Trans.”, “SO2”, and “SO3” represent the Point Cloud Aggregation Encoder Block, translation invariance, SO(2) equivariance, and SO(3) equivariance, respectively. The results show that, compared to the first row which corresponds to the original DP3 setup, incorporating the Aggregation Encoder Block together with the canonical representation effectively facilitates the learning of 3D policies.
In the point cloud encoder ablation study (Table III), we compared three encoders: the original MLP-based DP3 encoder, PointNet++ [46], and a graph-based encoder DGCNN [49]. These encoders were evaluated under different policies, including DP3, EquiBot, CP-SO3, and CP-SO2. To explain EquiBot’s limited performance on RoboMimic tasks, we also included its original encoder, the Vector Neuron SO(3)-equivariant encoder. Across all policies, when using an encoder other than the DP3 encoder, the success rate in MimicGen tasks dropped to nearly zero. Similarly, in the Push T task, the success rate remained notably limited.
@@
Finally, to examine canonical policy’s compatibility with different action heads, we compared diffusion model and flow matching [15] on Stack D1 and Push T. The results are presented in Table IV. It is noteworthy that diffusion and flow matching exhibit task-specific preferences. For instance, in Stack D1, diffusion outperforms flow matching for nearly all policies, whereas in Push T, flow matching leads to better results. Despite these differences, canonical policy consistently outperforms the other baselines under both types of action heads.
TABLE II: Task success rates (%) in the ablation study across different configurations.
@@
$\checkmark$ $\checkmark$ $\times$ $\checkmark$ $68\pm 4$ $37\pm 5$ $41\pm 4$ $68\pm 5$ $40\pm 7$ $85\pm 2$
$\checkmark$ $\checkmark$ $\checkmark$ $\times$ $76\pm 5$ $44\pm 7$ $38\pm 6$ $67\pm 8$ $52\pm 8$ $87\pm 3$
@@
TABLE III: Task success rates (%) for different encoders across policies and tasks.
@@
PointNet++ $0\pm 0$ $0\pm 0$ $0\pm 0$ $1\pm 1$ $0\pm 0$ $13\pm 1$
DGCNN $0\pm 1$ $0\pm 0$ $0\pm 0$ $0\pm 0$ $0\pm 0$ $22\pm 1$
CP-SO3 DP3 Encoder $50\pm 5$ $36\pm 6$ $29\pm 2$ $57\pm 8$ $39\pm 5$ $82\pm 3$
PointNet++ $0\pm 0$ $0\pm 0$ $0\pm 0$ $1\pm 1$ $0\pm 1$ $28\pm 3$
DGCNN $0\pm 0$ $0\pm 0$ $0\pm 0$ $0\pm 1$ $0\pm 0$ $17\pm 1$
CP-SO2 DP3 Encoder $57\pm 3$ $37\pm 4$ $30\pm 4$ $67\pm 6$ $51\pm 4$ $84\pm 2$
PointNet++ $0\pm 1$ $0\pm 0$ $0\pm 0$ $3\pm 3$ $1\pm 1$ $32\pm 1$
DGCNN $0\pm 0$ $0\pm 0$ $0\pm 0$ $0\pm 0$ $0\pm 0$ $27\pm 1$
@@
TABLE IV: Comparison of success rates (%) between Diffusion and Flow Matching on Stack D1 and Push T.

### v2

@@ V Simulation Experiments
In this section, we provide a comprehensive evaluation of canonical policy in simulation environments. We first present an experiment to validate the $\mathrm{SE}(3)$ equivariance property of canonical policy. We then conduct benchmark comparisons with several representative point cloud-based baselines across a diverse set of manipulation tasks. We further perform ablation studies to highlight the contributions of key components in our method. Finally, we present a dedicated analysis of the reliability of the rotation estimator introduced in Section IV-A, examining its performance under noisy inputs and illustrating the results with feature visualizations.
@@ V-A Experiment Setup
DP3 [DP3] is a diffusion-based policy that directly processes 3D point clouds using a simple MLP encoder.
iDP3 [iDP3] builds upon DP3 by replacing the MLP encoder with a multi-layer 1D convolutional encoder. This architecture is designed to extract multi-resolution features from point clouds, capturing more detailed and hierarchical point cloud information.
EquiBot [Equibot] is a $\mathrm{SIM}(3)$-equivariant policy that uses the same equivariant encoder as canonical policy, built on the Vector Neuron framework [VN]. It modifies the diffusion U-Net backbone using equivariant linear, pooling, and normalization layers. To conform with the Vector Neuron framework, EquiBot reformulates low-dimensional states and actions into vector and scalar representations.
Figure 6: Visualization of the 12 simulated manipulation tasks used in the simulation benchmark. Subfigures (a)–(h) are from MimicGen [robomimic], (i) is from Diffusion Policy [diffusion_policy], and (j)–(l) are from EquiBot [Equibot]. In each subfigure, the left image shows the initial state, and the right image shows the goal state of the task.
@@
MimicGen Tasks [robomimic] are visualized in Fig. 6(a)–(h). These tasks, such as coffee making and table organization, simulate realistic household manipulation scenarios. Point cloud observations are obtained from four camera views. To ensure computational efficiency without compromising performance, all point clouds are uniformly downsampled to 256 points using farthest point sampling [pointnet++].
Push T Task is shown in Fig. 6(i). Originally a 2D task from Diffusion Policy [diffusion_policy], we adapt it to 3D by extending each of the 9 keypoints with a zero z-coordinate, forming a sparse 3D point cloud. The goal is to push a T-shaped block to a placed T-shaped target.
Bimanual Tasks are depicted in Fig. 6(j)–(l). These dual-arm manipulation tasks are adapted from the released EquiBot dataset [Equibot], which provides 1024 point clouds as observation.
@@
In canonical policy, we employ absolute control for the MimicGen and Push T tasks, and relative control for the bimanual tasks. The difference between the two control modes lies in the choice of action space used during the SE(3) inverse and forward transformations. See Section IV-B for more details.
@@ V-B Validation of SE(3) Equivariance
Figure 7: Canonical pose visualization across tasks. Seen and novel scenes are aligned to a shared canonical frame via estimated SE(3) inverse transforms.
Figure 8: Quantitative evaluation of different policies across tasks. Curves show success rates over training epochs for seen scenes and novel scenes, with and without random $\mathrm{SE}(3)$ transformations. Canonical Policy maintains stable performance under both seen and novel conditions, while baseline policies degrade under novel conditions.
@@
As illustrated in Fig. 7, we visualize three MimicGen tasks: Stack D1, Hammer Cleanup D1, and Coffee D2, to demonstrate the SE(3)-equivariance property of canonical policy. As discussed in Section III-C, the left column of Fig. 7 presents point cloud observations from the training dataset, which are considered seen scenes and denoted as $\mathcal{D}$. Next, random SE(3) transformations are applied to generate novel scenes $\mathcal{D}^{\mathrm{rd}}$, shown in the middle column. Finally, by applying data canonicalization pipeline described in Section IV-A, we first center each transformed point cloud by subtracting its mean, then estimate the associated rotation using an $\mathrm{SO}(3)$-equivariant network (Equation (8)), and finally apply the inverse rotation to each point cloud, aligning it to a shared canonical frame, as visualized in the right column of Fig. 7.
To quantitatively assess the $\mathrm{SE}(3)$ equivariance of Canonical Policy, we design an experiment under four conditions, as shown in Fig. 8: (i) seen scenes (training set; solid red), (ii) seen scenes with random $\mathrm{SE}(3)$ transforms (solid blue), (iii) novel scenes (evaluation set; dashed red), and (iv) novel scenes with random $\mathrm{SE}(3)$ transforms (dashed blue). All policies are trained on the training set and tested under all four conditions.
For the Canonical Policy, as described in Section IV-B, both the observation and the low-dimensional inputs are mapped to a canonical frame using the estimated transformation; actions are predicted in this frame and then mapped back using the inverse transform. Consequently, samples that differ only by $\mathrm{SE}(3)$ transforms but belong to the same equivariant group share the same canonical representation. Hence, as shown in the last column of Fig. 8, training on seen data and testing on either seen or novel data, with or without random $\mathrm{SE}(3)$ transforms, yields similar success rate curves.
In contrast, baseline policies that do not enforce $\mathrm{SE}(3)$ equivariance generalize poorly: performance drops on novel scenes and degrades sharply, often to near zero, when random $\mathrm{SE}(3)$ transforms are introduced at evaluation. These findings verify the theoretical motivation in Section IV: canonicalization removes pose variability before policy inference, leading to robustness across arbitrary rigid transforms.
@@ V-C Benchmark with Point Cloud Policies
As discussed in Section IV-A, we provide two versions of canonical policy: one with translation invariance and full $\mathrm{SO}(3)$ equivariance, referred to as CP-$\mathrm{SO}3$; and another with translation invariance and in-plane $\mathrm{SO}(2)$ equivariance, referred to as CP-$\mathrm{SO}2$. We will use these abbreviations throughout the following experimental section.
Table II summarizes the performance of various policies across 12 tasks. Specifically, for the 8 MimicGen tasks, we generate 200 demonstrations [equidiff] and train each method for 250 epochs. For the Push T task, which also contains 200 demonstrations [diffusion_policy], each method is trained for 1000 epochs. For the bimanual tasks, each with 50 demonstrations [Equibot], training is conducted for 2000 epochs per method.
All experiments are conducted using three random environment seeds: 42, 43, and 44. For each seed, we evaluate the policies over the last 10 epochs and report the average. For MimicGen and Push T, each rollout is conducted across 50 different environment initializations. For the three bimanual tasks, each rollout consists of 10 single trials. Thus, the mean score reported in Table II for each task and each method is the average over 30 evaluation results. The best result for each task is shown in bold, and the second-best is underlined.
As shown in Table II, both CP-SO3 and CP-SO2 outperform the baseline point cloud policies (DP3, iDP3, and EquiBot) on all 8 MimicGen tasks and the 3D Push T task. Notably, CP-SO2 achieves the highest success rates in 6 out of the 9 single-arm tasks, supporting our hypothesis that $\mathrm{SO}(2)$ equivariance is particularly well-suited for scenarios where the $z$-axis remains consistently upward. In the three bimanual tasks, all policies perform similarly in Object Covering and Box Closing, achieving success rates close to 100%. However, for Cloth Folding, CP-SO3 and CP-SO2 achieve the second and third best performances, respectively, following EquiBot.
@@
Interestingly, while EquiBot achieves the best performance across all three bimanual tasks and ranks third on the 3D Push T task, it performs poorly on the MimicGen tasks. We hypothesize that this discrepancy arises from the point cloud encoder used in EquiBot [DP3], which will further explored in our ablation study on point cloud encoders.
TABLE II: Task success rates (%) across different policies and tasks.
@@
iDP3 $40\pm 6$ $45\pm 7$ $76\pm 1$ $1\pm 6 99$$9\pm 1$ $100\pm 0$
EquiBot $1\pm 1$ $5\pm 2$ $77\pm 4$ $8\pm 2 99$$9\pm 1$ $100\pm 0$
[…]
We report the mean success rate, computed as the average over the final 10 checkpoints, each evaluated in 50 environment initializations. Results are further averaged over 3 training seeds, yielding 150 evaluations per policy. Bold indicates the best performance; underline denotes the second-best. “CP-SO3” refers to canonical policy with SO(3) rotation equivariance, and “CP-SO2” denotes the variant with SO(2) equivariance. All policies are evaluated on point cloud inputs with only 3D coordinates, without using color information.
@@ V-D Ablation Study
We conducted comprehensive ablation studies focusing on five aspects: (i) the contributions of different modules within the canonical policy, (ii) the effects of various point cloud encoders, (iii) the role of scaling in task performance via comparison of $\mathrm{SE}(3)$ and $\mathrm{SIM}(3)$, (iv) the impact of key parameters in the $\mathrm{SO}(3)$-equivariant network on estimation performance, and (v) the influence of different action heads. In all experiments, success rates are computed as the mean over the final 10 checkpoints, each evaluated under 50 environment initializations with random seed 42.
As shown in Table III, the abbreviations “Aggre.”, “Trans.”, “SO2”, and “SO3” represent the Point Cloud Aggregation Encoder Block, translation invariance, SO(2) equivariance, and SO(3) equivariance, respectively. The results show that, compared to the first row which corresponds to the original DP3 setup, incorporating the Aggregation Encoder Block together with the canonical representation effectively facilitates the learning of 3D policies.
In the point cloud encoder ablation study (Table IV), we compared three encoders: the original MLP-based DP3 encoder, PointNet++ [pointnet++], and a graph-based encoder DGCNN [DGCNN]. These encoders were evaluated under different policies, including DP3, EquiBot, CP-SO3, and CP-SO2. To explain EquiBot’s limited performance on RoboMimic tasks, we also included its original encoder, the Vector Neuron SO(3)-equivariant encoder. Across all policies, when using an encoder other than the DP3 encoder, the success rate in MimicGen tasks dropped to nearly zero. Similarly, in the Push T task, the success rate remained notably limited.
@@
Furthermore, EquiBot adopts $\mathrm{SIM}(3)$ equivariance, which extends $\mathrm{SE}(3)$ equivariance by additionally scaling the entire scene to improve robustness to variations in object size. To evaluate this effect within our canonical policy, we conduct an ablation study comparing policies with and without scaling across four MimicGen tasks, as shown in Table V. Results are mixed: for both CP-SO3 and CP-SO2, scaling improves performance on two tasks but worsens it on the other two, leading us to adopt $\mathrm{SE}(3)$ canonicalization in the final model.
Next, we analyze two key hyperparameters of the $\mathrm{SO}(3)$-equivariant network in Fig. 3 that mainly determine the parameter budget of the entire policy: the number of repeat layers and the feature dimension of the equivariant linear layers. We evaluate their impact on the Stack D1 task for both CP-SO3 and CP-SO2, with results summarized in Table VI. Table VI(a) reports the proportion of parameters allocated to the $\mathrm{SO}(3)$-equivariant branch relative to the full policy (including the $\mathrm{SO}(3)$-equivariant branch, the point cloud encoder, and the diffusion action head). Two main observations emerge. First, the $\mathrm{SO}(3)$-equivariant branch is lightweight: under our default configuration (4 repeat layers, 48 feature dimensions) it accounts for only $\sim$3.2% of total parameters. Second, feature dimension has a stronger effect on parameter count than repeat depth, yielding a greater monotonic increase as the dimension grows. Performance trends in Table VI(b) (CP-SO3) and Table VI(c) (CP-SO2) further show that success rates improve with larger branch capacity, consistent with the intuition that a higher-capacity equivariant estimator is more robust to noisy inputs. This will be discussed in detail in the next subsection.
Finally, to examine canonical policy’s compatibility with different action heads, we compared diffusion model and flow matching [flow_matching] on Stack D1 and Push T. The results are presented in Table VII. It is noteworthy that diffusion and flow matching exhibit task-specific preferences. For instance, in Stack D1, diffusion outperforms flow matching for nearly all policies, whereas in Push T, flow matching leads to better results. Despite these differences, canonical policy consistently outperforms the other baselines under both types of action heads.
TABLE III: Task success rates (%) in the ablation study across different configurations.
@@
[…]
@@
TABLE IV: Task success rates (%) for different encoders across policies and tasks.
@@
PointNet++ $\pm 0 0±0$$\pm 0$ $0\pm 0$ $1\pm 1$ $0\pm 0$ $13\pm 1$
DGCNN $\pm 1 0±0$$\pm 0$ $\pm 0 0±0$$\pm 0$ $0\pm 0$ $22\pm 1$
[…]
@@
TABLE V: Ablation of scaling ($\mathrm{SIM}(3)$) versus no scaling ($\mathrm{SE}(3)$) on canonical policy performance across tasks.
Scaling Stack D1 Mug. D1 Ham. D1 Coffee D2
CP-SO3 $\times$ $68\pm 4$ $37\pm 5$ $68\pm 5$ $40\pm 7$
$80\pm 5$ $41\pm 5$ $65\pm 6$ $39\pm 3$
CP-SO2 $\times$ $76\pm 5$ $44\pm 7$ $67\pm 8$ $52\pm 8$
$69\pm 4$ $47\pm 7$ $71\pm 5$ $47\pm 5$
TABLE VI: Ablation on two key hyperparameters of the $\mathrm{SO}(3)$-equivariant network and performance on Stack D1.
Feature Dimensions Repeat Layers Average
1 2 4
12 0.20% 0.21% 0.22% 0.21%
24 0.80% 0.82% 0.87% 0.83%
48 3.11% 3.20% 3.39% 3.23%
(a) Proportion of $\mathrm{SO}(3)$-equivariant branch parameters relative to total policy parameters.
Feature Dimensions Repeat Layers Average
1 2 4
12 $68\!\pm\!3$ $62\!\pm\!4$ $61\!\pm\!4$ $64\!\pm\!5$
24 $73\!\pm\!5$ $63\!\pm\!4$ $77\!\pm\!4$ $71\!\pm\!7$
48 $74\!\pm\!4$ $70\!\pm\!4$ $68\!\pm\!4$ $70\!\pm\!5$
(b) Success rate ($\%$) of CP-SO3 with varying branch size.
Feature Dimensions Repeat Layers Average
1 2 4
12 $75\!\pm\!4$ $81\!\pm\!5$ $70\!\pm\!4$ $75\!\pm\!6$
24 $78\!\pm\!4$ $79\!\pm\!4$ $81\!\pm\!3$ $79\!\pm\!4$
48 $85\!\pm\!3$ $87\!\pm\!4$ $76\!\pm\!5$ $83\!\pm\!6$
(c) Success rate ($\%$) of CP-SO2 with varying branch size.
TABLE VII: Comparison of success rates (%) between Diffusion and Flow Matching on Stack D1 and Push T.